This question asks:

Calculate the change in entropy that occurs in the system when 1.00 mol of methanol (CH3OH) vaporizes from a liquid to a gas at its boiling point (64.6 degrees Celcius). For methanol, enthalpy of vaporization = 35.21 kJ/mol

I'm getting the correct answer (104) but I'm getting a negative for some reason, and the book says it's positive, but the formula changeinEntropy = -enthalpy of system/T, so why isn't this number negative?
Answer #1

Formula given,

Change in entropy for system = enthalpy of system/T

But for surrounding, Change in entropy = -enthalpy of system/T

Enthalpy of vaporization = 35.21 kJ/mol = 35210 J/mol

Temperature (T) = 64.6oC = 64.6 + 273.15 = 337.75 K

So, Change in entropy for system = (35210 J /mol)/337.75 K = 104 J/mol.K

so, It will be positive for system but negative for surrounding